OpenStudy (anonymous):

Whats the f' and f'' of : f(x)= (x^2)/9 + (x-4)^(2/3)?

OpenStudy (comm.dan):

Was it given to you?

OpenStudy (anonymous):

No, this is why I ask the question haha

OpenStudy (comm.dan):

Ok

OpenStudy (comm.dan):

Do you have an idea of how to solve functions?

OpenStudy (anonymous):

Yeah but I have some difficulties with that one! Can you help me or not?

OpenStudy (agent0smith):

\[\Large f(x) = \frac{ 1 }{ 9 }x^2 + (x-4)^{2/3}\] yes?

OpenStudy (anonymous):

yes !

OpenStudy (agent0smith):

Do you know how to find the derivative of the 1/9 x^2? Use the power rule: multiply by the exponent, reduce the exponent by one.

OpenStudy (haseeb96):

this is the correct answer

OpenStudy (anonymous):

\[\frac{ 2((x-4)^{4/3}-1) }{ 9(x-4)^{4/3} }\]

OpenStudy (anonymous):

this is the answer but I can't arrive to this answer by myself and I don't know why so that's why I need to know the complete calcul :(

OpenStudy (agent0smith):

Don't worry about the answer yet. Do it step by step.

OpenStudy (anonymous):

I do this but it doesn't work..

OpenStudy (agent0smith):

The answer will come from algebra, first you have to do calculus... do you know how to differentiate each term?

OpenStudy (anonymous):

no :p I need the solution

OpenStudy (anonymous):

The function, the first derivative and the second derivative from left to right.\[\left\{\frac{x^2}{9}+(x-4)^{2/3},\frac{2 x}{9}+\frac{2}{3 \sqrt[3]{x-4}},\frac{2}{9}-\frac{2}{9 (x-4)^{4/3}}\right\} \]

OpenStudy (agent0smith):

\[\Large f(x) = \frac{ 1 }{ 9 }x^2 + (x-4)^{2/3}\] \[\Large f'(x) = 2*\frac{ 1 }{ 9 }x^{2-1}+ \frac{ 2 }{ 3 }(x-4)^{2/3 -1}\]

OpenStudy (lukecrayonz):

Do you know the rules of finding a derivative?

OpenStudy (anonymous):

yes I know the rules but I cant do this number

OpenStudy (anonymous):

and thanks for the people that try to help me :)

OpenStudy (haseeb96):

see it, i forgot the last step but now i have done it

OpenStudy (anonymous):

Okay, its hard to read it but thanks anyway haha :)